OpenStudy (anonymous):

12/4= /5 wat is the blank

OpenStudy (unklerhaukus):

15

OpenStudy (preetha):

UnkleRhaukus, it may help her if you explained how you arrived at that solution.

Directrix (directrix):

12/4 = x / 5 4x = 60 x = 15

OpenStudy (preetha):

And does that make sense Snowball? Are you able to do one on your own? Thanks Directrix.

OpenStudy (anonymous):

i think u multiply 12 and 5 and then get answer that doesn' tmake sence

OpenStudy (preetha):

Ok, I see. When you have an equation, you can do the same thing on both sides, it remains an equation. So if you add 10 to the right hand side, you can add 10 to the left hand side and it will still be an equation. With me?

OpenStudy (preetha):

Now, we are trying to find out what is the number that is above 5, so we represent the blank by a x.

OpenStudy (preetha):

The next step is to simplify it. We want all the terms with x on one side of the = and all the terms without x on the other side.

OpenStudy (msmr):

I would cross multiply.

OpenStudy (anonymous):

i am with msmer

OpenStudy (unklerhaukus):

\[12/4=3\] \[x/5=3\]\[x=15\]

OpenStudy (preetha):

Multiply both sides by 5. On the left you have, 5 x (12/4) and on the other side you have 5 x (x/4) What do you get then ? What happens when you multiply 5 into x/4

OpenStudy (msmr):

This means: \[\frac{12}{4} = \frac{x}{5}\] To cross multiply, you multiply the numerator of one fraction with the denominator of the other fraction.

OpenStudy (preetha):

I like Unkle's solution as well.

OpenStudy (msmr):

5*12 = 4*x

OpenStudy (unklerhaukus):

why do it the hard way for?

OpenStudy (msmr):

60 = 4x

OpenStudy (msmr):

x = 15

OpenStudy (anonymous):

GIVE ME THE EASY WAY ONE STEP AT A TIME OK DRAW IT

OpenStudy (preetha):

Ok, snowball, go back to my last post, Are you with me?

OpenStudy (anonymous):

no draw it

OpenStudy (msmr):

unklerhaukus: for this particular problem, your solution would be easier because 12/4 simplifies to 3. However, in a case such as 13/4 where it does not simplify into a nice number, cross multiplying will be much easier - I thought it would be good to show snowballhockey how to set that up.

OpenStudy (preetha):